OpenStudy (anonymous):

find the limit of x approaches 1 in cotx/(cscx+1) i know the answer should be around (1/2) but i dont know how to get it @wio

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

I would start by multiplying top and bottom by the conjugate, which is \(\csc x - 1\)

OpenStudy (anonymous):

this would get me ((cotx)(cscx+1))/((cscx-1)(cscx+1))

OpenStudy (anonymous):

what would i do from there?

OpenStudy (anonymous):

any luck solving the problem?

OpenStudy (anonymous):

Expand the bottom

OpenStudy (anonymous):

how would expanding the bottom help?

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

so i get csc^2x-1, where do i go from there

OpenStudy (anonymous):

There is a trig identity for \(csc\) and \(\cot\)

OpenStudy (anonymous):

which trig identity would it be?

OpenStudy (anonymous):

Wait, this is the limit as \(x\to 1\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

any luck?

OpenStudy (anonymous):

Okay \[ \csc^2 x-1= \cot^2x \]This is what I was trying to get you to realize.

OpenStudy (anonymous):

i got that part, so i got \[1/\cot ^2x\]

OpenStudy (anonymous):

where do i go from there?

OpenStudy (anonymous):

Well the \(\cot\) cancel out, learing only one \(\cot\) at the bottom.

OpenStudy (anonymous):

what would the limit be then? \[\lim_{x \rightarrow 1} 1/\cot x\]

OpenStudy (anonymous):

It should become: \[ \lim_{x\to 1}\frac{\csc x -1}{\cot x} \]

OpenStudy (anonymous):

\[ \lim_{x\to 1}\frac{\csc x -1}{\cot x}= \lim_{x\to 1} (\tan x \csc x -\tan x) \]

OpenStudy (anonymous):

\[ \lim_{x\to 1}\frac{\csc x -1}{\cot x}= \lim_{x\to 1} (\tan x \csc x -\tan x)= \lim_{x\to 1} (\sec x -\tan x) \]

OpenStudy (jhannybean):

Now you just take the limit. That should be pretty straightforward.